OpenStudy (anonymous):

Which set of side lengths could form a triangle similar to a triangle with side lengths 11, 14, and 22? 22, 25, 33 22, 28, 42 33, 42, 66 33, 44, 66

ganeshie8 (ganeshie8):

take first option, and test if all side lengths have same ratio :- 22, 25, 33 22/11 = 2 25/14 = 1.xx (not 2, so reject)

ganeshie8 (ganeshie8):

take second option, and test if all side lengths have same ratio :- 22, 28, 42 22/11 = 2 28/14 = 2 42/22 = 1.xx (not 2, so reject)

ganeshie8 (ganeshie8):

take third option, and test if all side lengths have same ratio :- 33, 42, 66 33/11 = 3 42/14 = 3 66/22 = 3 so all sides have equal ratio ! so this is your guy !